OpenStudy (anonymous):

Sam picked a card from a standard deck, looked at it, and then put it back. He then picked a second card. What is the probability that Sam picked a heart or a king on either pick? one over thirteen sixteen over fifty two seventeen over fifty two sixteen over fifty three

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

How many hearts and how many kings in a standard deck of cards?

OpenStudy (anonymous):

4 kings and 15 hearts

OpenStudy (anonymous):

*13 hearts

OpenStudy (anonymous):

So total how many hearts and kings?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

which is basically 17/52

OpenStudy (anonymous):

Think about 17 is that correct?

OpenStudy (anonymous):

I am not sure.

OpenStudy (anonymous):

i think there are 13 hearts and and another 3 kings for a total of \(16\) not \(17\) to pick from

OpenStudy (anonymous):

OK there are 13 hearts (A,2,3,4,5,6,7,8,9,10,JQ,K) that is correct and there are 4 kings (KS,KC,KD,KH) and that is correct....but have we counted the king of hearts twice?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

then it is a matter of how to compute "on either pick" meaning on pick one, pick two, or both

OpenStudy (anonymous):

for this i would use \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\]

OpenStudy (anonymous):

Would you mind helping me for this one?

OpenStudy (anonymous):

actually i don't see the answer i would get as one of the choices maybe i am doing it wrong

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

probability he gets it on the first pick is \(\frac{4}{13}\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and since he replaces the card, the probability he gets it on the second pick is also \(\frac{4}{13}\)

OpenStudy (anonymous):

Yeah I believe so.

OpenStudy (anonymous):

now you are asked the probability he gets it on either pick in math that should mean on the first pick, on the second pick or on both picks

OpenStudy (anonymous):

no the probability is 16/52

OpenStudy (anonymous):

13 hearts and 3 kings

OpenStudy (anonymous):

for that we would use \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] with \[P(A)=P(B)=\frac{4}{13}\] and \[P(A\cap B)=(\frac{4}{13})^2\]

OpenStudy (anonymous):

the last one because they are independent @nelsonjedi \(\frac{16}{52}=\frac{4}{13}\)

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

It is 16/53.

OpenStudy (anonymous):

ooh i see!! what a poorly worded question!

OpenStudy (anonymous):

when you are asked "on either pick" they do not mean when you pick twice!! they just mean what is it on a pick!

OpenStudy (anonymous):

Oh nevermind sorry I was looking at another response

OpenStudy (anonymous):

ok so it is \(\frac{4}{13}\) but the question was written strangely

OpenStudy (anonymous):

the real answer is \(\frac{88}{169}\) but they want \(\frac{16}{52}\) so give them that one

OpenStudy (anonymous):

A poll was conducted in states A, B, and C to determine which candidate voters would most likely vote for. The results are displayed in the table below: State A State B State C Total Liberal 20 21 12 53 Conservative 20 14 13 47 Total 40 35 25 100 What is P(liberal | State B)? 0.21 0.35 0.4 0.6

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Kris wanted to understand whether students at her school were in favor of an extended school day. She surveyed some students and displayed the results in the table below: In favor Opposed Undecided Grade 9 6 4 8 Grade 10 10 11 9 Grade 11 12 15 11 Grade 12 15 6 14 If the principal randomly selects a student in grade 10 from this survey, what is the probability that the student is opposed to extending the school day? 0.09 0.29 0.30 0.31

OpenStudy (anonymous):

Do you think you can help @kropot72

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kropot72):

@gabylovesu Are you sure that you have copied the table and the choices of answer correctly? The reason that I ask is that none of the choices appear to be correct.

OpenStudy (anonymous):

Yup I did

OpenStudy (kropot72):

The way I look at it is this. We are given that the randomly selected student is in grade 10. The required probability is therefore: \[\frac{grade\ 10\ number\ against}{total\ grade\ 10\ students}\]

OpenStudy (anonymous):

The function f(x) = 15(2)x represents the growth of a frog population every year in a remote swamp. Elizabeth wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Elizabeth's purposes? f(x) = 15(2 to the one half power)2x f(x) = 15(22)the x over 2 power f(x) = fifteen halves(2)x f(x) = 30(2)x @ganeshie8

OpenStudy (anonymous):

Izzy's heartbeat rate was measured on an EKG machine. After conversion, the function produced could be modeled by a sine function, and the wave produced a maximum of 8, minimum of −2, and period of ð. Which of the following functions could represent Izzy's EKG read-out? f(x) = 8 sin ðx − 2 f(x) = 5 sin ðx + 3 f(x) = 5 sin 2x + 3 f(x) = 2 sin 2x + ð

OpenStudy (anonymous):

@ganeshie8

OpenStudy (triciaal):

@kropot72 do you think if the student was randomly selected then should consider the ways of choosing the opposed 10th grade student from the total number surveyed I think this has combined probability grade and how voted 41 opposed/total # surveyed and 11/30 10th grade opposed

OpenStudy (triciaal):

@gabylovesu Kris wanted to understand whether students at her school were in favor of an extended school day. She surveyed some students and displayed the results in the table below: I think the answer should be 0.09 11- 10th grade opposed from 121 surveyed 11/121 = 0.09

OpenStudy (kropot72):

@triciaal It is known that the randomly selected student is in grade 10. What is the probability that a student in grade 10 is opposed? Surely the answer is the fraction (total number opposed in grade 10) divided by (total number in grade 10): \[\frac{11}{30}=0.367\] Lets look at the probability that a randomly selected student if opposed: \[P(opposed)=\frac{total\ opposed}{total\ students}=\frac{36}{121}=0.2975\] On the basis of the above, do you really think that 0.09 can be correct?

OpenStudy (triciaal):

yes because there are still 11 random in grade 10

OpenStudy (triciaal):

looking at 2 conditions grade 10 and opposed

OpenStudy (kropot72):

My understanding of the question is that the principal goes to grade 10 and randomly selects a student. Therefore only the students in grade 10 are to be considered in the calculation.

OpenStudy (triciaal):

the principal randomly selects a student in grade 10 grade 10 and opposed 30/121*11/36 = 0.07 very close to 0.09 rounding maybe

OpenStudy (triciaal):

English is difficult the agreed solution.

OpenStudy (kropot72):

It is not possible for there to be only a 9 in 100 chance for the student to be opposed.

OpenStudy (triciaal):

not just opposed!!! 2 conditions are met #1 10th grade and #2 opposed

OpenStudy (kropot72):

But the question already gives the information that the student is in grade 10. Therefore the probability that the principal selects a student who is in grade 10 does not need to be calculated.

OpenStudy (triciaal):

yes because he was randomly selected from the different grades that were in the survey the first condition. I am serious when I said English is difficult. Read again how it is written. I understand your position.

OpenStudy (triciaal):

by the way I wrote "he" that should be "the student" because it did not say if it was a boy or girl